sandra is 36 years old. her son is 1 years old. how many years ago was sandra 3 times older than her son?
Answer by Alan3354(69443)

sandra is 36 years old. her son is 1 years old. how many years ago was sandra 3 times older than her son?
------------
The difference is 35 years.
3 times older = 4 times as old
-------
m = mother's age
s = son's age
---
m = s+35
m = 4s
---
4s = s+35
s = 35/3 years
m = 140/3 years
------
Years ago = 36 - 140/3 = negative age
No solution.